# Resonance of an object floating within a surface wavefield

### v2

@@ Abstract
We examine the interaction between floating cylindrical objects and surface waves in the gravity regime. Since the impact of resonance phenomena associated with floating bodies, particularly at laboratory scales, remains underexplored, we focus on the influence of the floaters’ resonance frequency on wave emission. We study the response of floating rigid cylinders to external mechanical perturbations. Using an optical reconstruction technique to measure wave fields, we study the natural resonance frequency of floaters with different sizes and their response to incoming waves. The results indicate that the resonance frequency is influenced by the interplay between the cylinder geometry and the solid-to-fluid density ratio. Experiments with incoming waves demonstrate that floaters diffract waves, creating complex patterns of diffusion and retro-diffusion while radiating secondary waves that interfere with the incident wavefield. Minimal wave generation is observed at resonance frequencies. These findings can provide insights for elucidating the behavior of larger structures, such as sea ice floes, in natural wave fields.
@@ I INTRODUCTION
Wave-structure interactions [1, 2] are encountered in various engineering applications, such as naval architecture and offshore operations, including ships, platforms, and buoys [3, 4, 5]. Interactions between waves and floating objects also arise in the marginal ice zone (MIZ), the polar region partially covered by ice floes formed of frozen sea ice [6]. The extent of the MIZs is determined by numerous effects, including wave-floes interaction [7, 8, 9], which has been theoretically studied for deformable cylinders [10]. Experimental measurements have also been conducted in wave basin to rationalize wave-floe interactions and to model the scattering of surface waves by floes [11, 12, 13, 14, 15]. Numerical simulations further extend these efforts [16, 17, 18]. A range of geometries, from floating spheres [19, 20] to complex offshore structures [21, 22, 23, 24], have been investigated. Floating bodies with simple geometry, such as cylindrical objects, exhibit pronounced heaving, pitching, and surging motions when subjected to anisotropic wave fields [25, 26, 27, 28, 29]. However, to our knowledge, there are no quantitative measurements of the wave field generated by a floating object oscillating in an external wave field.
In the laboratory-scale experiment we present here, we study the emission of waves by partially immersed cylinders. In particular, we compare the wave field generated by a free oscillating cylinder to the wave field induced by its oscillations when an external wave field triggers them. When the frequency of the external wave field is close to the floater’s natural frequency, minimal wave re-emission is observed in the transverse direction.
The manuscript is organized as follows. In Section II, we experimentally characterize the resonance frequency of floating cylindrical bodies and propose a theoretical model. In Section III, we study the interaction of these bodies with an externally generated wave field, highlighting how the body’s natural frequency affects this interaction. In all cases, the system behavior is characterized using an optical technique to reconstruct the liquid surface [30, 31].
@@ II.1 Experiments
As floating objects, we use polyethylene cylinders with density $\rho=0.95\,\rm{g/cm^{3}}$. The cylinder radius $R$ ranges from $R=7.5\,\rm{mm}$ to $30.0\,\rm{mm}$, and the height $h$ ranges from $h=5.0\,\rm{mm}$ to $20.0\,\rm{mm}$ (Fig 1b). Cylinders float at the surface of water in a tank with dimensions $78\,\rm{cm}\times 38\,\rm{cm}\times 19\,\rm{cm}$. The water depth $H=5.0\,\rm{cm}$ ensures that the surface waves are in the deep water regime in the entire working frequency range. We investigate a relatively wide range of aspect ratios ($h/R$), from $0.17$ to $2.67$. Due to their thickness and Young’s modulus, the floaters can be considered rigid bodies in the working frequency range. To ensure reproducible wetting conditions and to prevent overwashing, we cover the external surface of each cylinder with adhesive tape.
@@
Figure 2: Wave field and resonance frequency of floating cylinders experiencing instantaneous vertical impacts. (a) Top view of a floating cylinder ($R=30\,\rm mm$, $h=15\,\rm mm$) displaced by a rod (as depicted in the setup of Fig. 1a) over a checkerboard pattern with square size $a_{c}=4.0\,\rm mm$. (b) Water surface topography at time $t=1.39\,\rm s$ post impact. (c) Fourier spectrum $|\langle\hat{\zeta}\rangle_{x,y}|(f)$ of the wave field generated by the object perturbation shown in (b), with the resonance frequency $f_{0}=3.16\,\rm Hz$ indicated in red. (d) Measured resonance frequency as a function of the floaters’ height $h$ and for different radii $R$. Each data point corresponds to an individual experiment.
@@
The wavefield is recorded from the instant of impact until the waves reflect from the tank boundaries. The reconstructed surface $\zeta(x,y,t)$ (see Fig. 2b) is analyzed using Fourier analysis. We first compute $\hat{\zeta}_{x,y}(f)$, the Fourier spectrum in time at all (spatial) locations of the 2D reconstructed wave field. We then average the Fourier spectra in space over the entire reconstructed area. Fig. 2c shows the absolute value of $<\hat{\zeta}_{x,y}(f)>$ as a function of frequency $f$ in logarithmic scale. For all our experiments, we observe one prominent peak with frequency in the range $2-6$ Hz and a few secondary peaks at much larger frequencies (typically for $f>15$ Hz). To accurately estimate the frequency of maximum emission $f_{0}$, we fit the peak with a parabolic function. We also extract the amplitude of the resonance peak from this interpolation. By recording the motion of the floater from the side, we checked that the frequency $f_{0}$ measured from the wave field corresponds to the oscillation frequency of the vertical motion of the cylinder. Figure 2d shows the frequency $f_{0}$ of maximum emission as a function of the cylinder thickness $h$ and for different radii. We observe that $f_{0}$ decreases with both the cylinder thickness and the cylinder radius. This indicates that both parameters play a role in determining the heaving resonance frequency $f_{0}$.
@@ II.2 Modeling
Historically, floating objects have been studied, most notably in naval contexts, where engineers sought to understand the motion of ships. Complex variable techniques have been exploited to calculate the added mass of two-dimensional hulls atop a fluid of infinite depth as early as 1929 [32]. An improved investigation of the heaving motion of two-dimensional bodies was performed in later studies [25, 33, 34, 35], which also consider finite-depth effects. The motion of three-dimensional floating bodies has also been investigated in the case of vertical circular cylinders [36] and spheres [37]. The referenced works rely mainly on expansion formulae and integral expressions in order to solve for the ensuing motion of floating obstacles. In the present work, we utilize similar theoretical notions, but we invoke a simplifying geometric approximation so that an analytical expression for the oscillation frequency observed in our experiments could be attained in terms of the geometrical properties of the disk. The approximation agrees well with our data (see Fig. 4).
We describe heave oscillations of a buoyant cylinder floating on the free surface of a water bath with density $\rho_{w}$. We denote the cylinder’s submerged height with $h_{s}$. (Fig. 3a). In the limit of high-frequency oscillations, we derive an expression for the added mass and deduce the natural frequency of oscillation observed in experiments.
@@
For a stationary, partially submerged cylinder, the equilibrium depth is given by the vertical force balance, $\boldsymbol{F}=\boldsymbol{0}$, where
$\boldsymbol{F}$ $=$ $\left(-\rho_{w}\pi R^{2}gz-\rho\pi R^{2}hg\right)\hat{\boldsymbol{z}},$ (1)
$h_{s}$ $=$ $\frac{\rho}{\rho_{w}}h.$ (2)
For a small perturbation $\delta z(t)$ from the equilibrium position, a naive force balance on the floating object gives $\rho\pi R^{2}h\ddot{\delta z}=-\rho_{w}\pi R^{2}g\delta z$, which yields the natural frequency of oscillation $\Omega=\sqrt{\rho_{w}g/(\rho h)}$. However, such a calculation overlooks two important considerations. First, the surrounding fluid must be accelerated along with the cylinder, increasing the effective inertia. Second, the waves at the free surface impose a reaction force on the cylinder.
@@
The boundary value problem for a heaving cylinder is inherently complex, even when assuming potential flow and linearized boundary conditions. In the following, we use a high-frequency approximation to derive an asymptotic expression for the resonant frequency. We take the limit $\Omega^{2}L/g\rightarrow\infty$, with $L$ the typical length of the floating body, $g$ the acceleration due to gravity, and $\Omega$ the oscillation frequency. The problem then simplifies as the boundary condition on the free surface reduces to $\phi=0$, where $\phi$ is the velocity potential, while the standard boundary conditions remain applicable elsewhere (see Bai [35] and Ursell [33, §2]). Note that for a typical experiment, with $L=2R$ and $\Omega=2\pi f_{0}$, giving $\Omega^{2}L/g\sim 1-4$. Though these values are not extremely large, we still take the high-frequency limit as the first approximation to our problem.
We use the high frequency limit to compute the added mass by reflecting the submerged portion of a body and then halving the resultant added mass, referred to as the method of duplicated models [38, p. 135] (see the sketch in figure 3(c)). For a vertical translation of the body at some velocity, $\phi(x,y,-z)=-\phi(x,y,z)$ by symmetry, which implies $\phi(x,y,0)=0$. As a consequence, the potential $\phi$ around the shape sketched in figure 3(c) restricted to the domain $z<0$ is a solution to the boundary value problem outlined by Bai [35]. Since the flow magnitude is symmetric about $z=0$, half the energy of the flow around the translating shape in figure 3(c) is in the upper half plane and half is in the lower half-plane. Thus, the added mass for the body in figure 3(b) is precisely half of the added mass for the shape in figure 3(c).
@@
To compute the added mass of a floating cylinder of radius $R$ and submerged portion $h_{s}$, we instead compute the added mass of an immersed cylinder of height $2h_{s}$ and then divide by a factor of two, according to the method presented in the previous section. To avoid matching issues at the cylinder corner, we approximate the cylinder by an ellipsoid with semi-axis $h_{s}$ along the vertical direction and radii $R$ along the other two directions. We compute the potential flow around an ellipsoid traveling along its axis at speed $U$ as the solution to a Laplace problem in ellipsoidal coordinates. Next, the kinetic energy of the flow is computed explicitly (see Lamb [), Eq. 11 of §114]lamb1945hydrodynamics. By considering the pure translational motion of the ellipsoid along its semi-axis of radius $h_{s}$, the translational added mass $M_{22}$ writes:
@@
$(\rho\pi hR^{2}+M_{22})\ddot{\delta z}+\rho_{w}\pi R^{2}g\delta z=0.$ (4)
@@
Note that, in this formulation, we have neglected the memory convolution term that becomes important as the wave field builds up [36, see Eq. 31], and thus the frequency is expected to be valid on short time scales or in situations in which the history integral remains small.
We compare the theoretical prediction with the cylinder resonance frequency $f_{0}$ measured experimentally in Fig 4. Our model is in good agreement with the experimental data. The agreement is robust for smaller values of $\Omega$, in agreement that our approximation is especially valid for low frequencies.
@@ III FLOATING OBJECT IN AN EXTERNAL WAVEFIELD
In the free-floating configuration, a cylinder is placed 10 cm away from the plate and farther from the lateral edges. We systematically vary the wave frequency $f$ while maintaining a constant plate displacement amplitude. Figs. 5c and 5d show the comparison between the wave field with and without the object. In the presence of the object, we observe the superposition of the incoming wave, the reflected waves, and the waves emitted by the oscillations of the cylinder. We focus on wave emission near the natural frequency $f_{0}$ of the floater.

### v3

@@ Abstract
We examine the interaction between floating cylindrical objects and surface waves in the gravity regime. Since the impact of resonance phenomena associated with floating bodies, particularly at laboratory scales, remains underexplored, we focus on the influence of the floats’ resonance frequency on wave emission. First, we study the response of floating rigid cylinders to external mechanical perturbations. Using an optical reconstruction technique to measure surface wave fields in both space and time, we study the natural resonance frequency of floats with different sizes. The results indicate that the resonance frequency is influenced by the interplay between the cylinder geometry and the solid-to-fluid density ratio. Second, these floating objects are placed in an incoming wave field. These experiments demonstrate that floats diffract incoming waves, while radiating secondary waves that interfere with the incident wavefield. Minimal wave generation is observed at resonance frequencies. These findings can provide insights for elucidating the behavior of larger structures, such as sea ice floes, in natural wave fields.
@@ I INTRODUCTION
Wave-structure interactions are encountered in various engineering applications, such as naval architecture or offshore operations, including ships, platforms and buoys. [1, 2]. An external wave field may induce motion of a floating body possessing up to six different degrees of freedom. Linear translation modes in the horizontal plane are called sway and surge, while vertical translation is referred to as heave. Angular translation modes are called roll, pitch and yaw [1]. A number of theoretical studies have aimed to better understand these motions and the corresponding pressure forces on floating structures. Ursell [3] investigated the vertical heave motion of an infinitely long circular cylinder placed horizontally at the water surface, and computed the 2D velocity potential describing cylinder oscillations in the case of a non viscous purely potential flow. This enabled the computation of the added mass – the effective inertia added to the cylinder due to fluid which is displaced with the object’s movement. Havelock [4] generalized these results in three dimensions by looking at the heaving motion of a floating half-immersed sphere. The added mass was computed as a function of the oscillation frequency. Havelock also computed the evolution of the damping coefficient with the motion frequency. Forces exerted by an incoming, monochromatic wave field on a fixed floating body were then studied by Newman [5]. Pressure forces have been computed for floating objects of various shapes including, for example, ellipsoids. Black [6] and Garett [7] were interested by the forces exerted on a vertical circular cylinder excited by a monochromatic wave. Yeung [8] then computed the added mass and damping coefficient for the heave, surge and pitching motion of the same object. All these theoretical studies brought precious knowledge to better understand the behaviour of floating structures in small amplitude wave fields.
The need for modelling polar climates and specifically wave interactions with sea ice motivated several studies on wave transmission by an array of floes. Indeed, sea ice is a brittle material that can be easily broken into fragments by surface waves, creating large areas of ice floes called Marginal Ice Zones [9]. These collections of ice floes are known to effectively attenuate incoming waves [10]. One of the major mechanisms that explains this decay is the wave scattering by individual ice floes which was studied theoretically by Wadhams [11]. Each ice floe is put into motion by the incoming wave, generating radiative wave fields which can interact and result in the general damping of the initial wave field. These models and observations motivated several experiments performed in wave tanks such as the works of Bennetts and Williams [12] or Yiew et al. [13]. Single floating discs were excited by a linear monochromatic wave. The disc motion was recorded using markers placed on the objects or accelerometers [12]. This enabled the tracking of the amplitudes of heave, surge and pitch motion as a function of the wave frequency and amplitude.
Recent interest for wave-energy converters also motivated several experimental studies. For instance, Kramer et al. [14] and Colling et al. [15] studied the heave motion of a spherical buoy during a free-decay drop test. The buoy was dropped from a small height so that the water surface disturbance remained linear. The spherical buoy then underwent damped harmonic oscillations. Colling et al. observed that the oscillation frequency as well as the attenuation rate both depend on the buoy draft. Using a multi camera system as well as wave gauges they were able to compare the motion of the buoy and the local water surface deformation to different numerical models. While these large scale experiments mostly focus on the oscillation of the floating body and only study the water deformation at some point of the surface, experiments at a smaller centimetric scale enable the reconstruction of water surface around an oscillating body in both space and time. Rhee et al [16] studied the waves generated by a centimeter-scale, vibrating floating robot thanks to Fast Checkerboard Demodulation (FCD) [17]. This method is a variation of the Synthetic Schlieren Imaging method as developed by Moisy et al. [18], now widely used for wave measurement at the laboratory scale [19, 20, 21]. Based on the difference of optical path length between a background pattern and the camera, these methods enable the computation of water surface distorsion.
In this paper we focus our attention on the wave field generated by the motion of a floating circular cylinder. Using a surface reconstruction technique, we resolve in space and time the wave field radiated by a circular centimetric disk when submitted to an impulsive perturbation. From this wave field, we measure the natural resonance frequency of the object. Using the same reconstruction method, we place the floe in an anisotropic monochromatic wave and record the new wave field which results from the combination of the incoming, radiated and diffracted fields. Close to the floater’s natural resonance frequency, we observe a minimal re-emission of waves in the direction transverse to the initial wave field. The manuscript is organized as follows. In Section II, we experimentally measure the resonance frequency of floating cylindrical bodies and propose a theoretical model. In Section III, we study the interaction of this bodies with an externally generated wave field, highlighting how the float’s natural frequency affects this external field.
@@ II.1 Experiments
As floating objects, we use polyethylene cylinders with density $\rho=0.95\,\rm{g/cm^{3}}$. The cylinder radius $R$ ranges from $R=7.5\,\rm{mm}$ to $30.0\,\rm{mm}$, and the height $h$ ranges from $h=5.0\,\rm{mm}$ to $20.0\,\rm{mm}$ (Fig 1b). Cylinders float at the surface of water in a tank with dimensions $78\,\rm{cm}\times 38\,\rm{cm}\times 19\,\rm{cm}$. The water depth $H=5.0\,\rm{cm}$ ensures that the surface waves are in the deep water regime in the entire working frequency range. We investigate a relatively wide range of aspect ratios ($h/R$), from $0.17$ to $2.67$. Due to their thickness and Young’s modulus, the floaters can be considered as rigid bodies in the working frequency range. To ensure reproducible wetting conditions and to prevent overwashing, we cover the external surface of each cylinder with adhesive tape.
@@
Figure 2: Wave field and resonance frequency of floating cylinders experiencing instantaneous vertical impacts. (a) Top view of a floating cylinder ($R=30\,\rm mm$, $h=15\,\rm mm$) displaced by a rod (as depicted in the setup of Fig. 1a) over a checkerboard pattern with square size $a_{c}=4.0\,\rm mm$. (b) Water surface topography at time $t=1.39\,\rm s$ post impact. (c) Fourier spectrum $|\langle\hat{\zeta}\rangle_{x,y}|(f)$ of the wave field generated by the object perturbation shown in (b). The inset in (c) depicts the same Fourier spectrum in log-log scale. The main peak represented by red dots is fitted with a Lorentzian function, plotted as a red line. (d) Measured resonance frequency as a function of the floaters’ height $h$ and for different radii $R$. Each data point corresponds to an individual experiment. The black dotted line represents the prediction $f_{0}=\frac{1}{2\pi}\sqrt{\frac{\rho_{w}g}{\rho h}}$ obtained without considering the floater added mass.
@@
The wavefield is recorded from the instant of impact until the waves reflect from the tank boundaries. The reconstructed surface $\zeta(x,y,t)$ (see Fig. 2b) is analyzed using Fourier analysis. We first compute $\hat{\zeta}_{x,y}(f)$, the Fourier spectrum in time at all (spatial) locations of the 2D reconstructed wave field. We then average the Fourier spectra in space over the entire reconstructed area. Fig. 2c shows the absolute value of $\hat{\zeta}_{x,y}(f)>$ as a function of frequency $f$ in logarithmic scale. For all our experiments, we observe one prominent peak with frequency in the range 262-6 Hz and a few secondary peaks at much larger frequencies (typically for $f>15$ Hz).
This prominent peak is normalized and approached by a Lorentzian function $\frac{1}{\sqrt{1+4\pi^{2}\left(\frac{f-f_{0}}{\mu}\right)^{2}}}$. From this fit, we extract with good precision the frequency of maximum emission $f_{0}$ as well as the damping coefficient $\mu$. This coefficient accounts for several dissipative phenomenons such as bulk and boundary viscous effects, capillary effects at the object walls and waves radiation. Values of this damping coefficient range between $2$ and $4\;\mathrm{rad.s^{-1}}$ resulting in damped oscillations. However, the frequency shift induced by this energy loss is small, resulting in a relative shift in the order of $2\;\%$ on the natural resonance frequency. We therefore consider that the frequency of maximum emission $f_{0}$ we measured is a good proxy of the natural resonance frequency of the floater.
Figure 2d shows the frequency $f_{0}$ of maximum emission as a function of the cylinder thickness $h$ and for different radii. Each point corresponds to a single experiment. For a single cylinder geometry, three different experiments are performed, resulting in a mean standard deviation of about $0.04\;\mathrm{Hz}$. We observe that $f_{0}$ decreases with both the cylinder thickness and the cylinder radius. This indicates that both parameters play a role in determining the heaving resonance frequency $f_{0}$.
@@ II.2 Modeling
Historically, floating objects have been studied, most notably in naval contexts, where engineers sought to understand the motion of ships. Complex variable techniques have been exploited to calculate the added mass of two-dimensional hulls atop a fluid of infinite depth as early as 1929 [22]. An improved investigation of the heaving motion of two-dimensional bodies was performed in later studies [3, 23, 24], which also consider finite-depth effects. The motion of three-dimensional floating bodies has also been investigated in the case of vertical circular cylinders [8] and spheres [4]. The referenced works rely mainly on expansion formulae and integral expressions in order to solve for the ensuing motion of floating obstacles.
In the present work, we utilize similar theoretical notions. In the limit of high-frequency oscillations, and using a simplifying geometric approximation, we derive an expression for the added mass and deduce the natural resonance frequency of oscillation of a cylinder. Theses approximations agree well with our data (see Fig. 4).
We describe heave oscillations of a buoyant cylinder floating on the free surface of a water bath with density $\rho_{w}$. We denote the cylinder’s submerged height with $h_{s}$. (Fig. 3a).
@@
We consider a vertical unit vector $\vec{e_{z}}$ oriented upward. At rest, the cylinder’s weight $\rho\pi R^{2}hg\vec{e_{z}}$ and the resultant of buoyancy forces $\rho_{w}\pi R^{2}h_{s}g\vec{e_{z}}$ compensate, resulting in the expression of the submerged depth $h_{s}=\frac{\rho}{\rho_{w}}h$
For a small perturbation $\delta z(t)$ from the equilibrium position, a naive force balance on the floating object gives $\rho\pi R^{2}h\partial_{tt}{\delta z}=-\rho_{w}\pi R^{2}g\delta z$, which yields the natural frequency of oscillation $\Omega=\sqrt{\rho_{w}g/(\rho h)}$. However, such a calculation overlooks two important considerations. First, the surrounding fluid must be accelerated along with the cylinder, increasing the effective inertia. Second, the waves at the free surface impose a reaction force on the cylinder.
@@
The boundary value problem for a heaving cylinder is inherently complex, even when assuming potential flow and linearized boundary conditions. In the following, we use a high-frequency approximation to derive an asymptotic expression for the resonant frequency. We take the limit $\Omega^{2}L/g\rightarrow\infty$, with $L$ the typical length of the floating body, $g$ the acceleration due to gravity, and $\Omega$ the oscillation frequency. The problem then simplifies as the boundary condition on the free surface reduces to $\phi=0$, where $\phi$ is the velocity potential, while the standard boundary conditions remain applicable elsewhere (see Bai [24] and Ursell [25, §2]). Note that for a typical experiment, with $L=2R$ and $\Omega=2\pi f_{0}$, giving $\Omega^{2}L/g\sim 1-4$. Though these values are not extremely large, we still take the high-frequency limit as the first approximation to our problem.
We use the high frequency limit to compute the added mass by reflecting the submerged portion of a body and then halving the resultant added mass, referred to as the method of duplicated models [26, p. 135] (see the sketch in figure 3(c)). For a vertical translation of the body at some velocity, $\phi(x,y,-z)=-\phi(x,y,z)$ by symmetry, which implies $\phi(x,y,0)=0$. As a consequence, the potential $\phi$ around the shape sketched in figure 3(c) restricted to the domain $z<0$ is a solution to the boundary value problem outlined by Bai [24]. Since the flow magnitude is symmetric about $z=0$, half the energy of the flow around the translating shape in figure 3(c) is in the upper half plane and half is in the lower half-plane. Thus, the added mass for the body in figure 3(b) is precisely half of the added mass for the shape in figure 3(c).
@@
To compute the added mass of a floating cylinder of radius $R$ and submerged portion $h_{s}$, we instead compute the added mass of an immersed cylinder of height $2h_{s}$ and then divide by a factor of two, according to the method presented in the previous section. To avoid complications arising at the cylinder corner, we approximate the cylinder by an ellipsoid with semi-axis $h_{s}$ along the vertical direction and radii $R$ along the other two directions. We compute the potential flow around an ellipsoid traveling along its axis at speed $U$ as the solution to a Laplace problem in ellipsoidal coordinates. Next, the kinetic energy of the flow is computed explicitly (see Lamb [27][Eq. 11 of §114]). By considering the pure translational motion of the ellipsoid along its semi-axis of radius $h_{s}$, the translational added mass $M_{22}$ becomes,
@@
$(\rho\pi hR^{2}+\frac{1}{2}M_{22})\ddot{\delta z}+\rho_{w}\pi R^{2}g\delta z=0.$ (2)
@@
Note that, in this formulation, we have neglected the memory convolution term that becomes important as the wave field builds up [28, see Eq. 31], and thus the frequency is expected to be valid on short time scales or in situations in which the history integral remains small.
We compare the theoretical prediction with the cylinder resonance frequency $f_{0}$ measured experimentally in Fig 4. Our model is in good agreement with the experimental data. The agreement is robust for smaller values of $\Omega$, obtained for the larger tested objects. We observe that the experimental results deviate from the theory for smaller radii, typically for a radius of R = 7.5 mm. This deviation may correspond to the fact that for elongated vertical cylinders, the assumption of pure heaving motion may break down.
@@ III FLOATING OBJECT IN AN EXTERNAL WAVEFIELD
In the free-floating configuration, a cylinder is placed 10 cm away from the wave generator and farther from the lateral edges. We systematically vary the wave frequency $f$ while maintaining a constant plate displacement amplitude. Figs. 5c and 5d show the comparison between the wave field with and without the object. In the presence of the object, we observe the superposition of the incoming wave, the reflected waves, and the waves emitted by the oscillations of the cylinder. We focus on wave emission near the natural frequency $f_{0}$ of the floater.